OpenStudy (anonymous):

when 4131 people attended a concert, the concert hall was 90% full. How many seats does the concert hall have ?

OpenStudy (anonymous):

4590

OpenStudy (rulnick):

capacity * 0.9 = 4131 so capacity = 4131 / 0.9 = ...

OpenStudy (anonymous):

4131= X*0.9 hence x= 4131/.9=4590

OpenStudy (luigi0210):

Make a proportion: \[\frac{ 4131 }{ x }=\frac{ 9 }{ 10 }\]